OpenStudy (anonymous):

what is 21.80 divided by 0.7 and cna u please show me the work ??

OpenStudy (anonymous):

\[21.8\div0.7 = 218\div7\]= 210/7 +8/7= 31+1/7

OpenStudy (anonymous):

\[\frac{21.80}{.7}=\frac{21.80}{.7}*\frac{10}{10}=\frac{218}{7}=\frac{217+1}{7}=\frac{217}{7}+\frac{1}{7}\]

OpenStudy (anonymous):

217/7=31 so answer is : 31+1/7

OpenStudy (anonymous):

so whats the answer in decimall ???

OpenStudy (anonymous):

31.1428571428etc...

OpenStudy (anonymous):

THANK YOU !! :)))))))))))))))))